1. The lifetimes of the light bulbs produced by Filip Company are normally distributed with a mean of 1,150 hours and a standard deviation of 175 hours.
  • What is the probability that a randomly chosen light bulb will have a lifetime of more than 1,000 hours?
  • What percentage of the light bulbs would be expected to last between 1,000 and 1,500 hours?
1
Expert's answer
2021-03-16T07:17:52-0400

What is the probability that a randomly chosen light bulb will have a lifetime of more than 1,000 hours?

"μ = 1150 \; hours \\

σ = 175 \;hours \\

P(X > 1000) = 1 -P(X<1000) \\

= 1 -P(Z< \frac{1000-1150}{175} \\

= 1 -P(Z<-0.8571) \\

= 1 -0.1957 \\

= 0.8043"

What percentage of the light bulbs would be expected to last between 1,000 and 1,500 hours?

"P(1000<X < 1500) = P(\frac{1000-1150}{175}<Z< \frac{1500-1150}{175}) \\

= P(-0.8571<Z<2) \\

= P(Z<2) -P(Z<-0.8571) \\

= 0.9772 -0.1957 \\

= 0.7815 \\

= 78.15 \%"